A company sells cashews for $7 per pound and almonds for $4 per pound. How many pounds of each should be combined to get 15 pounds of a mixture worth $5 per pound?
Answer by mananth(16946)

Price ---------------- quantity
Cashews $7.00 ---------------- x pounds
Almonds $4.00 ------ 15 - x pounds
Mixture $5.00 ---------------- 15
15
7 x + 4 ( 15 - x ) = 75
7 x + 60 - 4 x = 75
7 x - 4 x = 75 - -60
3 x = 15
/ 3
x = 5 pounds $7.00 Cashews
10 pounds $4.00 Almonds

m.ananth@hotmail.ca